How do I solve for X?
(x+2)/3 - (x - 3)/2=5/6

Answer by lekelly(20)

How do I solve for X?
(x+2)/3 - (x - 3)/2=5/6
Step 1: Find GCD, which is 6
2(x+2)/6 - 3(x-3)/6=5/6
Step 2: Simplify
2x+4/6 + -3x+9/6 = 5/6
Step 3: Rewrite
(2x+4+-3x+9)/6=5/6
Step 4: Regroup
((2x-3x)+(4+9))/6 = 5/6
Step 5: Simplify
(-x+13)/6=5/6
Step 6: multiply each side by 6 to get rid of denominator
-x+13=5
Step 7: Solve for x
-x=5-13
-x=-8
x=8 ANSWER